Can someone help me solve this problem?
I need to solve: 10x - 1/4 = 9x + 3/8
Thank you so much,
Lyn

Answer by funmath(2933)

Hi Lyn,
Solve: 10x - 1/4 = 9x + 3/8
10x-9x-1/4=9x-9x+3/8
x-1/4=3/8
x-1/4+1/4=3/8+1/4
x=3/8+2/8
x=(3+2)/8
x=5/8
:
Happy Calculating, Lyn!!!